1.in your own words, explain what problem - solving cognitive levels entails in relation to the skills to be demonstrated. Write down an example of an unseen, non routine problem. Illustrate how you would solve the problem.

2.chickens and sheep are kept in a pen. The total number of animals is 10. The total number of legs is 34.how many are chickens and how many are sheep?

3.at molebogeng station, a train arrives every 50 minutes. The first train stops at 7:00 a.m. How many trains have stopped at the station just before 11:00 p.m.?

4. The product of two prime numbers is 10 and their sum is 7.what are the two numbers?


1
Expert's answer
2021-08-23T07:57:03-0400

1

It involves understanding the problem and then use senses and experience to solve it.

unseen eg Vin bought 2 dozen packs of eggs for $2 a dozen. How much change did Van get back if he paid with a $20 bill? unseen question here is "How much did Van spend in all on the eggs?"


non routine problem e.g long division

  1. Understand
  2. Plan
  3. Execute
  4. Review


2

say there are p sheep's and c chickens

then p+c = 10. So, c = number of chickens = 20 - p

and they have 4p+2c = 34 legs

substitute 10-p for c:

"4p + 2(10-p) = 34\\

4p -2p = 34-20\\

2p = 14\\

p = 7\\

c = 10-7=3\\
c=3\\"


3

Firstly, we need to calculate the difference in the number of hours

From the question, we have 7 am to 11 pm

7 am to 7 pm is 12 hours

7 pm to 11 pm is 4 hours

So total is 16 hours

16 hours to minutes is multiplied by 60 minutes

We have this as;

16 "\times" 60 = 960 minutes

so to get the number of trains, we simply have to divide the number of minutes by 50

Mathematically, we have this as 960/50

= 19.2

Since we cannot have a fractional train stop, it means the number of trains that has stopped is 19


4

x+y=7

xy=10

x=7-y

y(7-y)=10

y=5,2

The two numbers are 5 and 2.

5 + 2 = 7

5 x 2 = 10